OpenStudy (anonymous):

Simplify: –2m3 • (3m)2 A. –6m2 B. –6m3 C. –18m2 D. –18m5

OpenStudy (mathstudent55):

This is your question, right? \(–2m^3 \cdot (3m)^2\)

OpenStudy (mathstudent55):

Use this rule: \( (ab)^m = a^mb^m \) for the \( (3m)^2 \) part first.